6. Find the absolute maximum and minimum values of fon D, where f(x,y)=x² – 2xy +2y and D={(x,y)|0SX 33,0 sys2).
Answer #1

- 2xy + 24 2х- 28 E 2x + 2 fray) f(x14) +7 (7) р: { (+}} ; 213, 04 Не} for maxima and minima, fx (@ry) = во - 24 - 29 - - #yAt Corner Corner point Thus AL (3,0) +12,7 ) = 9 At (012) fray) = At (3,2) (2,4)= 9-12+4=1 Af(0,0) f(1) = 0 and fmint -0 Thus